OpenStudy (anonymous):

4(x-7) = 5(x+2) x= What?

OpenStudy (anonymous):

x = x

jimthompson5910 (jim_thompson5910):

4(x-7)=5(x+2) 4x-28=5x+10 4x=5x+10+28 4x-5x=10+28 -x=10+28 -x=38 x=38/(-1) x=-38

OpenStudy (dumbcow):

distribute 4x -28 = 5x +10 -x = 38 x = -38

OpenStudy (anonymous):

\[4x-28=5x+10\] \[-28-10=x\] \[-38=x\]

OpenStudy (anonymous):

any more questions ?